OpenStudy (anonymous):

The solution x=1/2 is a solution to which of the following equations? 5x = 20 4= 15x -4x = 2 30 = 60x Is D the answer?

jimthompson5910 (jim_thompson5910):

Yes it is because plugging in x = 1/2 gives you 30 = 60x 30 = 60(1/2) 30 = 30 which is a true equation

OpenStudy (anonymous):

Thanks @jim_thompson5910 :)

jimthompson5910 (jim_thompson5910):

np